OpenStudy (anonymous):

simplify: (b^(n)*5^(n+1))/(5^(n)*b^(n+1))

OpenStudy (zzr0ck3r):

b^(n-(n+1))*5((n+1)-n)) = b^(-1)*5^1 = 5/b

OpenStudy (anonymous):

i Dont get it

OpenStudy (anonymous):

b^(n-(n+1)) * 5((n+1)-n)???

OpenStudy (zzr0ck3r):

x^a / x^b = x^(a-b)

OpenStudy (zzr0ck3r):

like 2^3/2^2 = 2^(3-2) = 2^1 = 2 or 3^5/3^6 = 3^(5-6) = 3^(-1) = 1/3^1 = 1/3

OpenStudy (anonymous):

so 5^((n+1)-n) not 5((n+1)-n)

OpenStudy (zzr0ck3r):

correct sorry

OpenStudy (zzr0ck3r):

b^(n-(n+1))*5^((n+1)-n)) = b^(-1)*5^1 = 5/b

OpenStudy (anonymous):

i was so confuse haha

OpenStudy (zzr0ck3r):

:) good you should have been:)

OpenStudy (anonymous):

hahaha

OpenStudy (zzr0ck3r):

I just did that to test you..keep you on your toes....jk

OpenStudy (anonymous):

thanks for your help

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

i have a question

OpenStudy (anonymous):

why is n-(n+1) =1

OpenStudy (anonymous):

no why is n-(n+1)=-1

OpenStudy (anonymous):

it should be 1 right? or what am I doing wrong??